# Learning parameter dependence for Fourier-based option pricing with tensor trains

### v5

@@ Abstract
A long-standing issue in mathematical finance is the speed-up of option pricing, especially for multi-asset options. A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor trains to compress high-dimensional tensors. Another usage of the tensor train is to compress functions, including their parameter dependence. Here, we propose a pricing method, where, by a tensor train learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters. As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices. We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ paths in terms of computational complexity while keeping comparable accuracy.
@@ I Introduction
Pricing options appropriately is needed for making a profit and managing the risk of loss in option trading. According to the theory in mathematical finance  [^1] As typical textbooks in this area, we refer to [1, 2]., the price of an option is given by the expectation of the discounted payoff in the contract with some stochastic model on the dynamics of underlying asset prices assumed. Except for limited cases with simple contract conditions and models, the analytical formula for the option price is not available and thus we need to resort to the numerical calculation. In the rapidly changing financial market, quick and accurate pricing is vital in option trading, but it is a challenging task, for which long-lasting research has been made. In particular, pricing multi-asset options, whose payoff depends on the prices of multiple underlying assets, is often demanding. Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
@@
Motivated by these points, recently, applications of quantum computing to option pricing are considered actively [^2] See [3] as a comprehensive review.. For example, many studies have focused on applications of the quantum algorithm for Monte Carlo integration [4], which provides the quadratic quantum speed-up over the classical counterpart. Unfortunately, running such a quantum algorithm requires fault-tolerant quantum computers, which may take decades to be developed.
@@
In light of this, applications of quantum-inspired algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9]  [^3] Tensor train is a mathematically equivalent to matrix product state (MPS) [10]. learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^4] See [13, 14] as reviews. and recently, it has also been utilized in various fields such as machine learning [15, 16, 17], quantum field theory [18, 19, 20], and partial differential equations [7, 21, 22, 23, 24]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [25, 26, 19, 27] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
Here, we would like to point out an issue in this TT-based method. That is, we need to rerun TCI to obtain TTs each time the parameters are changed. According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
@@
To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [28, 29, 30] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We use these tensor trains to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ up to $d=11$ (refer to FIG. 2). We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ II.1.2 Singular value decomposition
In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low. This is done by first canonicalizing the TT using QR decompositions, and then performing the compression via SVD, discarding singular values that are smaller than the threshold $\epsilon_{\mathrm{SVD}}$ set for each bond. For more technical details, readers are referred to Ref. [10, 9].
@@ III.3 Monte Carlo-based option pricing
where $\vec{x}_{1},\ldots,\vec{x}_{N_{\rm path}}$ are i.i.d. samples from $q(\vec{x}|\vec{x}_{0})$. On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., [32]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions. Besides, calculating the payoff $v$ with the normal variable $\vec{x}_{i}$ involves exponentiation. In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
@@ IV.1 Outline
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in Fig. 1. Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in Fig. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. We reduce the bond dimensions of these TTs using SVD. We then contract adjacent core tensors pairwise and obtain two tensor train operators (TTOs), also known as matrix product operator (MPO), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$. We contract these TTOs along the index vector $\vec{j}$ as follows:
@@ IV.2 Computational complexity
[…]
Here, we ignore the computational complexity of all the processes in FIG. 1 (a) and consider that in FIG. 1 (b) after we get the TT only. This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing. As discussed in Sec. VI, we can reasonably find such a situation in practice.
@@ V.1 Details
We do not have the exact price of the multivariate min-call option since there is no known analytic formula for it. Instead, we regard the option price computed by the MC-based method with very many paths, concretely $5\times 10^{7}$ paths, as the true value. The error of the option price computed by the proposed method is evaluated by its deviation from the true value. We take the MC-based method with $10^{5}$ paths, which is a typical number in option pricing in practical business, as the comparison target against our method. The average relative error and its standard deviation of the MC-based method with $10^{5}$ paths are calculated from its 20 runs. We assess the accuracy of our method by seeing if it falls within the statistical error range of the MC-based method.
An issue is that the number of possible combinations of the parameter $\vec{p}$ is $100^{d}$, and thus we cannot test all of them. Thus, we randomly select 100 combinations and perform option pricing for each of them. We compare the maximum among the relative errors of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
@@
TensorCrossInterpolation.jl for TCI [34] has not yet been made publicly available. Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft provided us with access to this. ITensors.jl [35] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [36]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
V.2 Results
@@
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational time of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. We set the tolerance of SVD for obtaining $\chi_{V}$, maximum bond dimension of $\tilde{V}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
V.2.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at less than 10. The details of this compression by SVD are described in Appendix A. The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of 6 or less, regardless of when $\eta$ is constant or $\eta$ is automatically determined for each asset. Therefore, the computational complexity is much lower than that of MC-based option pricing. The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
It has been observed that accuracy tends to decrease as $d$ increases. Consequently, increasing $d$ further might result in decreased accuracy, making it difficult to maintain accuracy within the MC-based method.

### v6

@@ Abstract
A long-standing issue in mathematical finance is the speed-up of option pricing, especially for multi-asset options. A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor trains to compress high-dimensional tensors. Another usage of the tensor train is to compress functions, including their parameter dependence. Here, we propose a pricing method, where, by a tensor train learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters. As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices. We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ and $10^{6}$ paths in terms of computational complexity while keeping comparable accuracy.
@@ I Introduction
Pricing options appropriately is needed for making a profit and managing the risk of loss in option trading. According to the theory in mathematical finance  [^1] As typical textbooks in this area, we refer to Refs. [1, 2]., the price of an option is given by the expectation of the discounted payoff in the contract with some stochastic model on the dynamics of underlying asset prices assumed. Except for limited cases with simple contract conditions and models, the analytical formula for the option price is not available and thus we need to resort to the numerical calculation. In the rapidly changing financial market, quick and accurate pricing is vital in option trading, but it is a challenging task, for which long-lasting research has been made. In particular, pricing multi-asset options, whose payoff depends on the prices of multiple underlying assets, is often demanding. Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
@@
Motivated by these points, recently, applications of quantum computing to option pricing are considered actively [^2] See Ref. [3] as a comprehensive review.. For example, many studies have focused on applications of the quantum algorithm for Monte Carlo integration [4], which provides the quadratic quantum speed-up over the classical counterpart. Unfortunately, running such a quantum algorithm requires fault-tolerant quantum computers, which may take decades to be developed.
@@
In light of this, applications of quantum-inspired classical algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present  [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9]  [^3] Tensor train is a mathematically equivalent to matrix product state (MPS) [10]. learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in Ref. [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^4] See [13, 14] as reviews. and recently, it has also been utilized in various fields such as machine learning [15, 16, 17], quantum field theory [18, 19, 20], and partial differential equations [7, 21, 22, 23, 24]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [25, 26, 19, 27] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
Here, we would like to point out an issue in this TT-based method. That is, we need to rerun TCI to obtain TTs each time the input parameters of option price are changed. According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
@@
To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [28, 29, 30, 31] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We use these tensor trains to evaluate the integral including parameter dependence and perform fast option pricing in response to various parameter changes (refer to FIG. 1). To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ and $10^{6}$ paths by a factor of $O(10^{4})$ and $O(10^{5})$, respectively (see Fig. 2). We also confirm numerically that in the tested cases, the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths, and, although worse in some cases, is overall comparable to that with $10^{6}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
In the context of TT-based approximations of high-dimensional functions, incorporating parameter dependence into TTs has been considered in some fields [28, 29, 30, 31]. However, to the best of our knowledge, our study is the first to take such an approach in FT-based option pricing in order to make it more efficient for varying parameters, which provides practical benefits in the rapidly changing financial market.
@@ II.1.2 Singular value decomposition
In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low. This is done by first canonicalizing the TT using QR decompositions, and then performing the compression via SVD, discarding singular values that are smaller than the tolerance $\epsilon_{\mathrm{SVD}}$ set for each bond. This tolerance is defined by
$\epsilon_{\mathrm{SVD}}=\frac{|\tilde{F}_{\mathrm{TT}}-\tilde{F}^{{}^{\prime}}_{\mathrm{TT}}|^{2}_{\mathrm{F}}}{|\tilde{F}_{\mathrm{TT}}|^{2}_{\mathrm{F}}},$ (4)
where $|\cdots|^{2}_{\mathrm{F}}$ indicates the Frobenius norm, $\tilde{F}_{\mathrm{TT}}$ is the TT obtained from TCI and $\tilde{F}^{{}^{\prime}}_{\mathrm{TT}}$ is the other TT after SVD. For more technical details, readers are referred to Ref. [10, 9].
@@ III.3 Monte Carlo-based option pricing
where $\vec{x}_{1},\ldots,\vec{x}_{N_{\rm path}}$ are i.i.d. samples from $q(\vec{x}|\vec{x}_{0})$. On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., Ref. [33]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions. Besides, calculating the payoff $v$ with the normal variable $\vec{x}_{i}$ involves exponentiation. In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
@@ IV.1 Outline
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in FIG. 1. Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in FIG. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. We reduce the bond dimensions of these TTs using SVD. We then contract adjacent core tensors pairwise and obtain two tensor train operators (TTOs), also known as matrix product operators (MPOs), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$. We contract these TTOs along the index vector $\vec{j}$ as follows:
@@ IV.2 Computational complexity
The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $\tilde{V}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{\tilde{V}}^{2})$ [10]. Here, we denote the maximum bond dimensions as $\chi_{V}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of all the processes in FIG. 1 (a) and consider that in FIG. 1 (b) after we get the TT only. This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing. As discussed in Sec. VII, we can reasonably find such a situation in practice.
@@ V.1 Details
We do not have the exact price of the multivariate min-call option since there is no known analytic formula for it. Instead, we regard the option price computed by the MC-based method with very many paths, concretely $5\times 10^{7}$ paths, as the true value. The error of the option price computed by the proposed method is evaluated by its deviation from the true value. We take the MC-based method with $10^{5}$ paths, which is a typical number in option pricing in practical business, as the comparison target against our method. We also run the MC method with $10^{6}$ paths, as the value when high accuracy is needed. As the error of the MC-based option price, we compute the half-width of its 95% confidence interval, namely $1.96\sigma/\sqrt{n_{\rm MC}}$, where $\sigma$ is the sample standard deviation of the discounted payoff in the MC run and the path number $n_{\rm MC}$ is $10^{5}$ or $10^{6}$. We assess the accuracy of our method by seeing if it falls within the statistical error range of the MC-based method.
An issue is that the number of possible combinations of the parameter $\vec{p}$ is $100^{d}$, and thus we cannot test all of them. Thus, we randomly select 100 combinations and perform option pricing for each of them. We compare the maximum among the errors of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
@@
TensorCrossInterpolation.jl [35] was used for learning tensor trains from functions appearing Fourier-based option pricing. ITensors.jl [36] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [37]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
VI Results
@@
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$. For the parameter set that yields this maximum error, we compute the half-width of the 95% confidence interval for the MC-based option price and denote it by $e_{\mathrm{MC},10^{5}}$ and $e_{\mathrm{MC},10^{6}}$ for the cases with $10^{5}$ and $10^{6}$ paths, respectively. The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{5}$ paths are denoted by $c_{\mathrm{MC},10^{5}}$ and $t_{\mathrm{MC},10^{5}}$, with those for $10^{6}$ paths being 10 times larger. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
VI.0.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in FIG. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10 when $\eta$ is constant. When $\eta$ is automatically determined for each asset, the bond dimensions are around 60 or 70. The details of this compression by SVD are described in Appendix. The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, regardless of when $\eta$ is constant or $\eta$ is automatically determined for each asset. Therefore, the computational complexity is much lower than that of MC-based option pricing.
We compare the TT-based results with the MC simulations with $10^{5}$ and $10^{6}$ paths. With $10^{5}$ paths, the TT-based option price is within the MC error range in all the tested cases. With $10^{6}$ paths, the TT-based method overall gives the errors comparable with the MC-based method, though it performes worse for $d=8$ and $d=9$.